Knowing that the minimum angle of deviation for the same prism is 30º, calculate the refractive index of the prism and the new angle of incidence at the first surface.

(3 marks)

1
Expert's answer
2022-01-13T09:21:47-0500

"\frac{n_2}{n_1}=\frac{\sin\alpha}{\sin\beta},"

"n_1=1,~\sin\alpha=1,"

"n_2=\frac 1{\sin\beta}=2.0."